Solve
a^3+4s^2-45a=0
Answers
A) 0,-5,9
B) 5,-9
C) -36,-9
D)0,5,-9
ps ^ means to the power of ....thanks

Found 2 solutions by Theo, ewatrrr:
Answer by Theo(13342)

the formula is:

a^3 + 4a^2 - 45a = 0

factor out an a to get:

a * (a^2 + 4a - 45) = 0

factor the a^2 + 4a - 45 to get:

a * (a + 9) * (a - 5) = 0

your solutions are:

a = 0
a = -9
a = 5

you confirm by substituting in your equation to see if it is true.

your equation is

a^3 + 4a^2 - 45a = 0

when a = 0, the equation becomes 0 + 0 - 0 = 0 which becomes 0 = 0

when a = -9, the equation becomes -729 + 324 + 405 = -729 + 729 = 0

when a = 5, the equation becomes 125 + 100 - 225 = 225 - 225 = 0

answer is confirmed as good.


Answer by ewatrrr(24785)


Hi
Assuming the s is an 'a'
a^3+4a^2-45a=0
solving
a(a^2 + 4a - 45) = 0
a(a + 9)(a - 5) = 0
a = 0
(a + 9)= 0
a = -9
(a - 5)= 0
a = 5
solutions are: 0, -9 and 5
answer d would be the correct choice